OpenStudy (anonymous):

what is the inverse of y=-3x+6/7

OpenStudy (anonymous):

first: x=-3y + 6/7 second: x - 6/7= -3y third: isolate y by dividing it by -3 like this… x - 6/7= -3y/-3 this cancels out the -3 on the right side. now just divid th left side of the equation by -3. x/-3 - 6/7 * -3 = y your answer will be f

OpenStudy (anonymous):

i meant f-1(x) = x/-3 + 18/7